Suppose the mean amount of cholesterol in eggs labeled “large” is 186 milligrams, with standard deviation 7 milligrams. Find the probability that the mean amount of cholesterol in a sample of 144 eggs will be within 2 milligrams of the population mean.
1
Expert's answer
2020-11-24T06:25:53-0500

μ"\overline{X}"=μ=186

σ"\overline{X}"=σ/"\sqrt{n}" =7/12=0.583

Z=(188−186)/0.583=3.43

P(184<"\overline{X}"<188)=P(-Z<x<Z)=P(-3.43<x<3.43)=0.9994